OpenStudy (anonymous):

Find the dimensions of a rectangle with area 343 m2 whose perimeter is as small as possible.

Vocaloid (vocaloid):

first take the square root of 343 to find the side length then multiply that by 4

Directrix (directrix):

Courtesy of @MikeyMaximum xy = 343 y = 343/x P(x) = Perimeter = 2x+2y P(x) = 2x + 686x^-1 dP/dx = 2 - 686/x^2 = 0 for maximum or minimum 2x^2 - 686 = 0 x = √ (343) = 7√7 y = 7√7 also Those are the dimensions. Concept from @Allan The minimum perimeter for a rectangle of any given area is a square whose side measure would be the square root of the area.

Directrix (directrix):

@hpacheco518